OpenStudy (anonymous):

The base of a cone has a radius of 6 cm. The height of the cone is 15 cm. What is the slant height of the cone? A. 4.58 B. 16.16 C. 19.21 D. 5.26

OpenStudy (anonymous):

a^2=b^2+c^2 B = The radius which is one side of the triangle. C =The slant which is the hypotenuse of the triangle

OpenStudy (anonymous):

B. 16.16

OpenStudy (anonymous):

a^2 + b^2 = c^2 since we're looking for the slant 6^2 + 15^2 = c^2

OpenStudy (anonymous):

yeah :P u got this

OpenStudy (anonymous):

thanks !